HELP WITH REQUIREMENT 3

Several years ago, Westmont Corporation developed a comprehensive budgeting system for planning and control purposes. While departmental supervisors have been happy with the system, the factory manager has expressed considerable dissatisfaction with the information being generated by the system. A report for the company's Assembly Department for the month of March follows: Assembly Department Cost Report For the Month Ended March 31 Actual Results Planning Budget Variances Machine-hours 15,000 20,000 Variable costs: Supplies $ 8,700 $ 9,300 $ 600 F Scrap 29,400 31,500 2,100 F Indirect materials 86,600 102,000 15,400 F Fixed costs: Wages and salaries 75,100 71,000 4,100 U Equipment depreciation 101,000 101,000 – Total cost $ 300,800 $ 314,800 $ 14,000 F After receiving a copy of this cost report, the supervisor of the Assembly Department stated, “These reports are super. It makes me feel really good to see how well things are going in my department. I can’t understand why those people upstairs complain so much about the reports.” For the last several years, the company’s marketing department has chronically failed to meet the sales goals expressed in the company’s monthly budgets. Required: 1. The company’s president is uneasy about the cost reports, identify at least two reasons. 2. What kind of reports should be used to give better insight into how well departmental supervisors are controlling costs? 3. Complete the new performance report for the quarter, based on Flexible Budget Performance approach. 4. Were costs well controlled in March?
Answer #1

1) Cost reports are ineffective since budgeted costs at one level of activity are compared to actual costs

at another level of activity.

2) Flexible Budget performance reports must be used

3) Cost Report

Particulrs Actual Spending Variances Flexible Budget Activity Variances Planning Budget
Machine Hours 15000 15000 20000
Variable Cost :-
Supplies 8700 1725 U 6975 2325 F 9300
Scrap 29400 5775 U 23625 7875 F 31500
Indirect Materials 86600 10100 U 76500 25500 F 102000
Fixed Costs:-
Wages and Salaries 75100 4100 U 71000 - 71000
Equipment Depreciation 101000 - 101000 - 101000
Total Cost 300800 21700 U 279100 35700 F 314800

Calculation of Flexible Budget of :-

Supplies = ($9300/20000)*15000 = $6975

Scrap = ($31500/20000)*15000 = $23625

Indirect Materials = ($102000/20000)*15000 = $76500